OpenStudy (anonymous):

What polynomial has roots of –1, –2, and 3 ?

OpenStudy (anonymous):

start with \[(x+1)(x+2)(x-3)\] and multiply out

jhonyy9 (jhonyy9):

calcule (x+1)(x+2)(x-3)

OpenStudy (anonymous):

x^3+2x^2-7x+2?? did i do that right?

jhonyy9 (jhonyy9):

(x2+3x+2)(x-3)=x3-3x2+3x2-9x+2x-6=x3-7x-6

OpenStudy (anonymous):

okay i seee thank u

jhonyy9 (jhonyy9):

yw

jhonyy9 (jhonyy9):

good luck bye